OpenStudy (anonymous):

An LCO Flat Screen TV at Best Buy increased in price from $900 to $1,200. What was the percent of increase?

OpenStudy (anonymous):

(1200-900) /900 = 0.33333 In percentage, 0.33333*100= 33%

OpenStudy (radar):

The increase in price was $300. The problem is to figure what percentage is 300 to 900. we can divide the 300 by 900 to get the decimal equivalent of the percentage 300/900=1/3=.333333333 Multiply that by 100 gives you a percentage of 33.33% increase in mark-up

OpenStudy (radar):

Lesson being be careful when shopping at Best Buy lol